OpenStudy (anonymous):

This Problem is worth 1 Point There are six chairs in a row. Person A picks one of the chairs at random. Then Person B picks a chair at random from the five chairs that remain. Find the chance that they pick chairs that are next to each other.

OpenStudy (anonymous):

one the chair has been selected, there are 5 chairs left and two next to the first chair chosen, making your answer \(\frac{2}{5}\)

OpenStudy (gorv):

ans will be 1/3 which is correct ... if p1 select 1st chair then p2 will select 2nd chair so probability= 1/6*1/5... now if p1 select any chair between 2 to 5 (probab=1/6)then p2 will have 2 choices either on the left or right(probab=2/5)...and probababilty for ist n last chair will be same so probability=1/6*1/5+1/6*2/5+1/6*2/5+1/6*2/5+1/6*2/5+1/6*1/5=1/3

OpenStudy (gorv):

@satellite73 sry bt ur ans is wrong